OpenStudy (anonymous):

Ally is analyzing a circle, y2 + x2 = 4, and a linear function g(x). Will they intersect? x g(x) 0 3 4 0 8 −3 Answers: Yes at positive x-coordinates or zero Yes at negative x-coordinates or zero Yes at negative and positive x-coordinates or zero No they will not instersect

OpenStudy (danjs):

put together the line equation y=mx+b from the points. put that in the circle equation for y see if it works (m*x + b)^2 + x^2 = 4 those will be x values of intersection(s) if any exist

OpenStudy (anonymous):

So whats the answer @DanJS

OpenStudy (danjs):

idk, try it out

OpenStudy (anonymous):

Im not sure how to do it... @DanJS

OpenStudy (danjs):

Line -- y=(-3/4)*x + 3 circle -- y^2 + x^2 = 4 substitution of line into circle [ (-3/4)*x +3 ] + x^2 = 4 solve, see if you get a value for x.